### v2

# A new hierarchy for complex plane curves

Abstract.
We define the type of a plane curve as the initial degree of the corresponding Bourbaki ideal. Then we show that this invariant behaves well with respect to the union of curves. The curves of type 00 are precisely the free curves and the curves of type $1$ are the plus-one generated curves. In this note we show first that line arrangements and conic-line arrangements can have all the theoretically possible types. In the second part we study the properties of the curves of type $2$ and construct families of line arrangements and conic-line arrangements of this type.
Key words and phrases:
Jacobian ideal, Jacobian module, exponents, free curve, plus-one generated curve, line arrangements, conic-line arrangements
2010 Mathematics Subject Classification:
Primary 14H50; Secondary 14B05, 13D02, 32S22
1. Introduction
Let $S=\mathbb{C}[x,y,z]$ be the polynomial ring in three variables $x,y,z$ with complex coefficients, and let $C:f=0$ be a reduced curve of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{2}$. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e. the homogeneous ideal in $S$ spanned by the partial derivatives $f_{x},f_{y},f_{z}$ of $f$, and by $M(f)=S/J_{f}$ the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$. Consider the graded $S$-module of Jacobian syzygies of $f$, namely
$D_{0}(f)=\{(a,b,c)\in S^{3}\ :\ af_{x}+bf_{y}+cf_{z}=0\}.$
According to Hilbert’s Syzygy Theorem, the graded Jacobian algebra $M(f)$ has a minimal resolution of the form
(1.1) $0\to F_{3}\to F_{2}\to F_{1}\to F_{0},$
where clearly $F_{0}=S$, $F_{1}=S^{3}(1-d)$ and the morphism $F_{1}\to F_{0}$ is given by
$(a,b,c)\mapsto af_{x}+bf_{y}+cf_{z}.$
With this notation, the graded $S$-module of Jacobian syzygies $D_{0}(f)$ has the following minimal resolution
$0\to F_{3}(d-1)\to F_{2}(d-1).$
We say that $C:f=0$ is an $m$-syzygy curve if the module $F_{2}$ has rank $m$. Then the module $D_{0}(f)$ is generated by $m$ homogeneous syzygies, say $r_{1},r_{2},...,r_{m}$, of degrees $d_{j}=\deg r_{j}$ ordered such that
$1\leq d_{1}\leq d_{2}\leq...\leq d_{m}.$
We call these degrees as the exponents of the curve $C$, and $r_{1},...,r_{m}$ as a minimal set of generators for the module $D_{0}(f)$. Note that $d_{1}=mdr(f)$ is the minimal degree of a non-trivial Jacobian relation in $D_{0}(f)$. It is known that
(1.2) $d_{1}\leq d_{2}\leq d_{3}\leq d-1,$
see for instance [17, Theorem 2.4]. On the other hand, for line arrangements we have a better bound, namely
(1.3) $d_{m}\leq d-2,$
see [26, Corollary 3.5], and for conic-line arrangement we have
(1.4) $d_{m}\leq d-1,$
see [9, Corollary 1.5].
Let $I_{f}$ denote the saturation of the Jacobian ideal $J_{f}$ with respect to the maximal ideal ${\bf m}=(x,y,z)$ in $S$. Consider the following local cohomology group, usually called the Jacobian module of $f$,
$N(f)=I_{f}/J_{f}=H^{0}_{\bf m}(M(f)).$
We set $n(f)_{k}=\dim N(f)_{k}$ for any integer $k$ and we introduce the freeness defect of the curve $C$ by the formula
$\nu(C)=\max_{j}\{n(f)_{j}\}$
as in [17]. Note that $C$ is free if and only if $N(f)=0$ and hence $\nu(C)=0$, and $C$ is nearly free if and only if $\nu(C)=1$, see for instance [6, 15] for the definitions of free and nearly free curves. Recall that the total Tjurina number of a given curve $C\subset\mathbb{P}^{2}$ is defined to be
$\tau(C)=\sum_{p\in{\rm Sing}(C)}\tau_{p},$
where ${\rm Sing}(C)$ denotes the set of all singular points of $C$ and $\tau_{p}$ is the Tjurina number of the singularity $(C,p)$.
Consider the general form of the minimal resolution for the Milnor algebra $M(f)$ of a curve $C:f=0$ that is assumed to be not free, namely
(1.5) $0\to\oplus_{i=1}^{m-2}S(-c_{i})\to\oplus_{j=1}^{m}S(1-d-d_{j})\to S^{3}(1-d)\to S,$
with $c_{1}\leq...\leq c_{m-2}$ and $d_{1}\leq...\leq d_{m}$. It follows from [20, Lemma 1.1] that one has
(1.6) $c_{j}=d+d_{j+2}-1+\varepsilon_{j},$
for $j=1,...,m-2$ and some integers $\varepsilon_{j}\geq 1$. Using [20, Formula (13)], it follows that one has
(1.7) $d_{1}+d_{2}=d-1+\sum_{i=1}^{m-2}\varepsilon_{j}.$
This formula implies the following result, for details see [17, Theorem 2.3].
Theorem 1.1.
Let $C:f=0$ be a reduced plane curve of degree $d$ and let $d_{1}$ and $d_{2}$ be the minimal degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f)$ as above. Then the following hold.
  1. (1)
    The curve $C$ is free if and only if $d_{1}+d_{2}=d-1$.
  2. (2)
    The curve $C$ is plus-one generated if and only if $d_{1}+d_{2}=d$.
  3. (3)
    The curve $C$ is neither free nor plus-one generated if and only if $d_{1}+d_{2}>d$.
For the definition of plus-one generated curves, see [1, 17]. In view of this result, it seems natural to introduce the following invariant, which leads to an interesting hierarchy among all plane curves.
Definition 1.2.
Let $C:f=0$ be a reduced plane curve of degree $d$ and let $d_{1}$ and $d_{2}$ be the two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f)$ as above. Then we say that $C$ has type $t(C)$, where the positive integer $t(C)$ is defined by the formula
$t(C)=d_{1}+d_{2}+1-d.$
The fact that this invariant plays a key role in the study of plane curves follows from the following alternative definition of it. Let $C:f=0$ be a reduced curve in $\mathbb{P}^{2}$ and $r_{1}$ be a minimal degree generator of the graded $S$-module $D_{0}(f)$. Then the corresponding Bourbaki ideal $B(C,r_{1})\subset S$ of $D_{0}(f)$ was introduced in [18, Section 5], and [17, Section 3], after being considered in a non-explicit way by du Plessis–Wall in [19] and the first author in [6]. Its properties were further explored in the recent paper [22]. It follows from the definition of the Bourbaki ideal $B(C,r_{1})$ that the type of the curve $C$ is just the initial degree of the homogeneous ideal $B(C,r_{1})$, namely
(1.8) $t(C)=\operatorname{indeg}B(C,r_{1})=\min\{s\in\mathbb{Z}\ :\ B(C,r_{1})_{s}\neq 0\}.$
Note that a curve $C$ is free exactly when $t(C)=0$, and a curve $C$ is plus-one generated exactly when $t(C)=1$. The type of a curve is a positive integer bounded by $22d-$ 2 in view of (1.2), and this upper bound is attained for instance by a smooth curve. For line arrangements the corresponding upper bound is $42d-$ 4 in view of (1.3), see Corollaries 1.9 and 1.11 below for more on this subject.
Our first result says that the type behaves well with respect to the union of two curves. Consider two reduced curves $C_{1}:f_{1}=0$ and $C_{2}:f_{2}=0$, with no common irreducible component. Then one has the following easy result, concerning the union curve $C=C_{1}\cup C_{2}:f=f_{1}\cdot f_{2}=0$, see [14, Theorem 5.1].
Proposition 1.3.
With the above notation, one has
$\max\{mdr(f_{1}),mdr(f_{2})\}\leq mdr(f)\leq\min\{mdr(f_{1})+\deg C_{2},mdr(f_{2})+\deg(C_{1})\}.$
The type of curves behaves similarly with respect to the union of curves, namely we have the following result.
Theorem 1.4.
Let $C_{1}:f_{1}=0$ and $C_{2}:f_{2}=0$ be two curves in $\mathbb{P}^{2}$, with no irreducible component in common. Then the union curve $C=C_{1}\cup C_{2}:f=f_{1}f_{2}=0$ satisfies
$t(C)\leq\min\{t(C_{1})+\deg(C_{2}),t(C_{2})+\deg(C_{1})\}.$
If $C_{1}$ is free and $C_{2}$ is a line, then this result is already known, see [8, 23].
Corollary 1.5.
If $C$ is a free curve and $Q$ is a conic intersecting $C$ in a finite number of points, then the union $C^{\prime}=C\cup Q$ is a curve of type at most $2$. Similarly, if $C$ is a plus-one generated curve and $L$ is a line, which is not an irreducible component of $C$, then the union $C^{\prime}=C\cup L$ is a curve of type at most $2$.
Remark 1.6.
Consider the Fermat curve $C_{1}:f_{1}=x^{d}+y^{d}+z^{d}=0$ and the line arrangement $C_{2}:f_{2}=x^{d}+y^{d}=0$. Then the curve $C_{1}$ has type $d-1$ since the exponents of $C_{1}$ are $d_{1}=d_{2}=d_{3}=d-1$. On the other hand, the union $C=C_{1}\cup C_{2}$ has type 0, since $C$ is a free curve, see [11, Corollary 1.6]. It follows that there is no interesting lower bound for the type of $C=C_{1}\cup C_{2}$, without involving more information on $C_{1}$ and $C_{2}$.
The next result tells us what happens for the type $t(C)$, in most cases, when the curve $C_{2}$ is generic, namely smooth and transversal to $C_{1}$.
Theorem 1.7.
Let $C_{1}:f_{1}=0$ be a curve and let $d_{1}\leq d_{2}$ be the two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f_{1})$. Let $C_{2}:f_{2}=0$ be a smooth curve, such that the intersection $C_{1}\cap C_{2}$ consists of $\deg C_{1}\deg C_{2}$ points. If
$d_{2}\leq\deg C_{1}-2,$
then the union curve $C=C_{1}\cup C_{2}:f=f_{1}f_{2}=0$ satisfies
$t(C)=t(C_{1})+\deg C_{2}.$
More precisely, if $d_{1}^{\prime}\leq d_{2}^{\prime}$ denote the first two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f)$, then
$d_{1}^{\prime}=d_{1}+\deg C_{2}\text{ and }d_{2}^{\prime}=d_{2}+\deg(C_{2})\leq\deg(C)-2.$
Example 1.8.
The condition $d_{2}\leq\deg(C_{1})-2$ is necessary in Theorem 1.7 as the following example shows. Let $C_{1}:f_{1}=x^{2}+y^{2}+z^{2}=0$ and $C_{2}:f_{2}=x=0$. Then $d_{1}=d_{2}=1$ and hence $t(C_{1})=1$. A direct computation shows that $C=C_{1}\cup C_{2}$ is a 3-syzygy curve with exponents $(1,2,2)$ and hence of type
$21+2-2=1<t(C_{1})+\deg(C_{2})=$ 2 .
Corollary 1.9.
Let ${\mathcal{A}}$ be a line arrangement in $\mathbb{P}^{2}$ containing at least 3 lines. Then one has
$30\leq t({\mathcal{A}})\leq\deg({\mathcal{A}})-$ 3 .
Conversely, for any integers $t$ and $d$ such that $30\leq t\leq d-$ 3, there are line arrangements in $\mathbb{P}^{2}$ such that $t({\mathcal{A}})=t$ and $\deg({\mathcal{A}})=d$.
To prove a result similar to Corollary 1.9 for conic-line arrangements is more complicated. The smooth conic $C_{1}$ in Example 1.8 satisfies $t(C_{1})=d-1$, while the conic-line arrangement $C$ from the same example satisfies $t(C)=d-2$. It is easy to check that 2 smooth conics meeting in 4 points give rise to conic-line arrangement ${\mathcal{C}}$ with $t({\mathcal{C}})=2=d-2$, since $d=\deg({\mathcal{C}})=4$. We believe that such examples are exceptional, and ask the following question.
Question 1.10.
Is it possible to list all the conic-line arrangements ${\mathcal{C}}$ satisfying
$t({\mathcal{C}})\geq deg({\mathcal{C}})-2\ ?$
On the positive side, we have the following analog of Corollary 1.9 for conic-line satisfying an additional condition.
Corollary 1.11.
Let ${\mathcal{C}}:f=0$ be a conic-line arrangement and let $d_{1}\leq d_{2}$ be the two smallest degrees of a minimal system of generators for the module of Jacobian syzygies $D_{0}(f_{1})$. If $d_{2}\leq deg({\mathcal{C}})-2$, then
$30\leq t({\mathcal{C}})\leq\deg({\mathcal{C}})-$ 3 .
Conversely, for any integers $t$ and $d$ such that $30\leq t\leq d-$ 3, there are conic-line arrangements ${\mathcal{C}}$ containing at least one conic in $\mathbb{P}^{2}$ such that $t({\mathcal{C}})=t$ and $\deg({\mathcal{C}})=d$.
In particular, for any degree $d\geq 3$, there are free conic-line arrangements ${\mathcal{C}}$ containing at least one conic such that $\deg({\mathcal{C}})=d$.
Remark 1.12.
The type of a line arrangement is not determined by the combinatorics. Indeed, the pair of arrangements
${\mathcal{A}}:f=xy(x-y-z)(x-y+z)(2x+y-2z)(x+3y-3z)(3x+2y+3z)$
$(x+5y+5z)(7x-4y-z)=0,$
and respectively by
${\mathcal{A}}^{\prime}:f^{\prime}=xy(4x-5y-5z)(x-y+z)(16x+13y-20z)(x+3y-3z)(3x+2y+3z)$
$(x+5y+5z)(7x-4y-z)=0,$
constructed by Ziegler [32] and Yuzvinski [31], have the same combinatorics and as exponents $(5,6,6,6)$ and respectively $(6,6,6,6,6,6)$. It follows that the type of ${\mathcal{A}}$ is 3, and the type of ${\mathcal{A}}^{\prime}$ is 4.
One can make a loose analogy of the type of curves introduced in Definition 1.2 with Arnold’s notion of modality of function germs. In the case of function germs, the 00-modal (alias the simple singularities), the $1$-modal (alias the unimodal singularities) and the $2$-modal singularities are classified, see for instance [2], and the complexity of the corresponding geometry, e.g. reflected in their Milnor lattices, is increasing as the modality increases.
Following this path, from now on in this note on we focus on curves $C$ having type $2$. The exponents and the corresponding Tjurina numbers of the curves of type $2$ are described in the following result, where the notation from (1.7) is used.
Proposition 1.13.
Let $C$ be a curve of type $2$. Then one of the following two cases occurs.
  1. (1)
    $C$ is a 3-syzygy curve and $\varepsilon_{1}=2$. If $(d_{1},d_{2},d_{3})$ are the exponents of $C$, then
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-2d_{1}-2d_{2}-2d_{3}.$
  2. (2)
    $C$ is a 4-syzygy curve and $\varepsilon_{1}=\varepsilon_{2}=1$. If $(d_{1},d_{2},d_{3},d_{4})$ are the exponents of $C$, then
    $\tau(C)=d_{1}^{2}+d_{1}d_{2}+d_{2}^{2}-2d_{1}-2d_{2}-d_{3}-d_{4}+1.$
Conversely, if $m=3$ and $\varepsilon_{1}=2$ or $m=4$ and $\varepsilon_{1}=\varepsilon_{2}=1$, then $t(C)=2$.
In case $(1)$ (resp. $(2)$), we say that $C$ is a curve of type $2A$ (resp. $2B$). For the curves $C$ of type 2, we also have simple formulas for the freeness defect $\nu(C)$, as shown below.
Proposition 1.14.
Suppose that $C$ has type $2A$ and exponents $(d_{1},d_{2},d_{3})$. Then the following hold.
  1. (1)
    If $d_{1}<d_{2}-2$, then $\nu(C)=2(d_{3}-d_{2})+4\geq 4.$
  2. (2)
    If $d_{1}=d_{2}-2$, then $\nu(C)=2(d_{3}-d_{1})\geq 4.$
  3. (3)
    If $d_{1}=d_{2}-1$, then $\nu(C)=2(d_{3}-d_{1})+2\geq 4.$
  4. (4)
    If $d_{1}=d_{2}$, then $\nu(C)=2(d_{3}-d_{2})+3\geq 3.$
Proposition 1.15.
Suppose that $C$ has type $2B$ and exponents $(d_{1},d_{2},d_{3},d_{4})$. Then the following hold.
  1. (1)
    If $d_{1}<d_{2}-2$, then $\nu(C)=d_{3}+d_{4}-2d_{2}+3\geq 3.$
  2. (2)
    If $d_{1}=d_{2}-2$, then $\nu(C)=d_{3}+d_{4}-2d_{1}-1\geq 3.$
  3. (3)
    If $d_{1}=d_{2}-1$, then $\nu(C)=d_{3}+d_{4}-2d_{1}+1\geq 3.$
  4. (4)
    If $d_{1}=d_{2}$, then $\nu(C)=d_{3}+d_{4}-2d_{2}+2\geq 2.$
The equality in Proposition 1.15 (4) holds if and only if $d_{1}=d_{2}=d_{3}=d_{4}$, see also [17, Theorem 3.11 and Example 3.12] in relation with such curves.
Let us look at the first examples of curves $C$ having type $2$.
Example 1.16.
(i) Consider the Bolza curve $C:x^{5}-y^{2}z^{3}-xz^{4}=0$. It is well-known that $C$ has a unique singularity of type $E_{8}$ and one can check that $C$ has type $2A$ with exponents $(2,4,4)$.
(ii) Consider a conic-line arrangement $\mathcal{CL}_{1}\subset\mathbb{P}^{2}_{\mathbb{C}}$ given by
$f_{1}=(x^{2}+y^{2}-z^{2})(y-z)(x^{2}-z^{2})=0.$
This arrangement is known to be free and it has $3$ singularities of type $A_{1}$ and $3$ singularities of type $A_{3}$, so this gives us $\tau(\mathcal{CL}_{1})=12$, see [24, Example 2.9]. If we add the conic $Q\,:y^{2}-xz=0$ to $\mathcal{CL}_{1}$, then the resulting arrangement $\mathcal{CL}$ has quasi-homogeneous singularities and has type $2B$ with exponents $(4,4,5,5)$.
(iii) Consider a conic arrangement $\mathcal{C}_{1}\subset\mathbb{P}^{2}_{\mathbb{C}}$ given by
$f_{1}=(-3x^{2}+xy+yz+xz)(-3y^{2}+xy+yz+zx)(-3z^{2}+xy+yz+zx)=0.$
This arrangement is known to be free with exponents $(2,3)$ and it has $3$ singularities of type $A_{5}$ and just one ordinary triple point $D_{4}$. If we add the conic $Q:f_{2}=x^{2}+y^{2}+z^{2}=0$, we get new conic arrangement $\mathcal{C}$ which has type $2A$ and exponents $(4,5,6)$. Note that the conic $Q$ is generic with respect to the arrangement $\mathcal{C}_{1}$, namely the intersection $\mathcal{C}_{1}\cap Q$ consists of $12$ points, which give rise to $12$ nodes for the arrangement $\mathcal{C}$
Remark 1.17.
One very natural way to obtain line arrangements of type $2$ is the following. Start with a free line arrangement ${\mathcal{A}}$ and let ${\mathcal{A}}_{i,j}$ be the line arrangement obtained from ${\mathcal{A}}$ by deleting two lines, namely $L_{i}$ and $L_{j}$. Then [4, Theorem 1.8] gives sufficient condition such that $t({\mathcal{A}}_{i,j})=2$. More precisely, with the notation from this result, we have the following possibilities. The arrangements in [4, Theorem 1.8] (1) have type $2B$, and the same holds for the arrangements in case (2), when $c_{1}<c_{2}$. On the other hands, the arrangements in [4, Theorem 1.8] (2) have type $2A$ if $c_{1}=c_{2}$ and $t({\mathcal{A}}_{i,j})\neq 1$. Moreover, in [4, Example 4.5], the author produces a free line arrangement ${\mathcal{A}}$ of $11$ lines, such that the arrangement ${\mathcal{A}}_{3,4}$ has type $2A$ and the arrangements ${\mathcal{A}}_{2,11}$ and ${\mathcal{A}}_{3,5}$ have type $2B$.
It is natural to ask the question whether in Corollary 1.5, in both cases, the two types $2A$ and $2B$ are possible for the curve $C^{\prime}$. Our next result gives a rather general construction to start with a plus-one generated line arrangement ${\mathcal{A}}^{\prime}$ and get arrangements ${\mathcal{A}}$ of type $2B$, by adding a generic line.
Theorem 1.18.
Let ${\mathcal{A}}^{\prime}$ be a line arrangement of type 1 with exponents $(m_{1},m_{2},m_{3})$ with $m_{1}\geq 3$. Let $L$ be a line such that the intersection $L\cap{\mathcal{A}}^{\prime}$ consists of $\deg{\mathcal{A}}^{\prime}=m_{1}+m_{2}$ points. Then the line arrangement ${\mathcal{A}}={\mathcal{A}}^{\prime}\cup L$ has type $2B$ and the exponents of ${\mathcal{A}}$ are
$(m_{1}+1,m_{2}+1,m_{3}+1,m_{1}+m_{2}-1).$
In the next result, we consider a line arrangement as a special type of conic-line arrangement, i.e. a conic-line arrangement where the number of conics is zero. The example of conic-line arrangement constructed in Theorem 1.19 is very different from the conic-line arrangements ${\mathcal{C}}$ constructed in the proof of Corollary 1.11, since here all the singularities situated on the conic are quasi-homogeneous, which is not the case for the singularity $({\mathcal{C}},(0:0:1))$ in general, see Remark 3.3.
Theorem 1.19.
Let ${\mathcal{C}}^{\prime}$ be a free conic-line arrangement with exponents $(m_{1},m_{2})$ with $m_{1}\geq 2$. Let $Q$ be a smooth conic such that the intersection $Q\cap{\mathcal{C}}^{\prime}$ consists of $2\deg{\mathcal{C}}^{\prime}=2(m_{1}+m_{2}+1)$ points. Then the arrangement ${\mathcal{C}}={\mathcal{C}}^{\prime}\cup Q$ has type $2A$ with exponents $(m_{1}+2,m_{2}+2,m_{1}+m_{2}+1)$.
Example 1.20.
An illustration of Theorem 1.19 occurs in Example 1.16 (iii), where $(m_{1},m_{2})=(2,3)$. Note that in Example 1.16 (ii) the condition that the intersection $Q\cap{\mathcal{A}}^{\prime}$ consists of $102\deg{\mathcal{A}}^{\prime}=2(m_{1}+m_{2}+1)=$ 10 points is not fulfilled, since $\tau(\mathcal{CL})=24$ by Proposition 1.13, and hence
$\tau(\mathcal{CL}_{1})+10=22\neq\tau(\mathcal{CL}).$
This explains why the type $2B$ may occur here.
Remark 1.21.
It is interesting to note that the formula for $\tau(C)$ in Proposition 1.13 (1) can be obtained by setting $d_{4}=d_{3}+1$ in formula for $\tau(C)$ in Proposition 1.13 (2). Similarly, all the formulas for $\nu(C)$ in Proposition 1.14 can be obtained by setting $d_{4}=d_{3}+1$ in the corresponding formulas for $\nu(C)$ in Proposition 1.15. In particular, the sets of exponents $(n_{1}+1,n_{2}+1,n_{1}+n_{2}-2)$ and $(n_{1}+1,n_{2}+1,n_{1}+n_{2}-2,n_{1}+n_{2}-1)$ yield the same Tjurina number $\tau(C)$ and the same freeness defect $\nu(C)$. For this reason the Tjurina number and the freeness defect cannot be used to show that the type $2B$ does not occur in Theorem 1.19. The same situation occurs below in Theorem 1.24. To discard the case $2B$ in both Theorems we need to use deep results from [27].
We have seen that Theorem 1.18 produces line arrangements of type $2B$. To get families of line arrangements of type $2A$, one can use the following construction. We start by constructing first a family of line arrangements ${\mathcal{A}}^{\prime}$ of type $1$ by taking the union of two pencils of lines. The following result may be well-known to the specialists.
Proposition 1.22.
Let ${\mathcal{B}}_{1}$ (resp. ${\mathcal{B}}_{2}$) be the arrangement consisting of $n_{1}$ lines passing through a point $a_{1}\in\mathbb{P}^{2}$ (resp. $n_{2}$ lines passing through a point $a_{2}\in\mathbb{P}^{2}\setminus{\mathcal{B}}_{1}$) and such that $1\leq n_{1}\leq n_{2}$ and $n_{1}+n_{2}>2$. Then the arrangement ${\mathcal{A}}^{\prime}={\mathcal{B}}_{1}\cup{\mathcal{B}}_{2}$ has type 00 and exponents $(1,n_{1}+n_{2}-2)$ for $n_{1}=1$, and type $1$ and exponents $(n_{1},n_{2},n_{1}+n_{2}-2)$ for $n_{1}\geq 2$.
In view of Theorem 1.18, this directly implies the following.
Corollary 1.23.
If we add a generic line to the line arrangement ${\mathcal{A}}^{\prime}$ from Proposition 1.22 with $n_{1}\geq 3$, then we get a line arrangement of type $2B$ with exponents
$(n_{1}+1,n_{2}+1,n_{1}+n_{2}-1,n_{1}+n_{2}-1).$
The next result tells us what happens when we add a line $L$ to the arrangement ${\mathcal{A}}^{\prime}$ in Proposition 1.22, passing through exactly one double point of ${\mathcal{A}}^{\prime}$.
Theorem 1.24.
Let ${\mathcal{A}}^{\prime}$ be the line arrangement from Proposition 1.22 and assume that $3\leq n_{1}\leq n_{2}$. Let $L$ be a line such that the intersection $L\cap{\mathcal{A}}^{\prime}$ consists of $n_{1}+n_{2}-1$ points. Then the arrangement ${\mathcal{A}}={\mathcal{A}}^{\prime}\cup L$ has has type $2A$ with exponents of $(n_{1}+1,n_{2}+1,n_{1}+n_{2}-2)$.
Finally we present an enumerative description of line arrangements of type $2$. Let us first fix the notation. Let $\mathcal{L}\subset\mathbb{P}^{2}$ be an arrangement of $d$ lines. We denote by $t_{k}:=t_{k}(\mathcal{L})$ the number of $k$-fold intersection points of $\mathcal{L}$ and we define the maximal multiplicity of $\mathcal{L}$ as follows:
[…]
Theorem 1.25.
Let $\mathcal{L}\subset\mathbb{P}^{2}$ be an arrangement of $d$ lines of type $2$. Then one has
$m(\mathcal{L})\geq\bigg{\lceil}\dfrac{4d}{d+5}\bigg{\rceil}.$
Moreover, the following hold.
  1. (1)
    If $\mathcal{L}$ has type $2A$ and exponents $(d_{1},d_{2},d_{3})$, then
    $\sum_{r\geq 2}(r-1)t_{r}\geq d_{1}d_{2}+2.$
  2. (2)
    If $\mathcal{L}$ has type $2B$ and exponents $(d_{1},d_{2},d_{3},d_{4})$, then
    $\sum_{r\geq 2}(r-1)t_{r}\geq d_{1}d_{2}+1.$
Both lower bounds for $\sum_{r\geq 2}(r-1)t_{r}$ for line arrangements of type $2$ given above are sharp and this follows from Example 1.26 below.
Example 1.26.
We start by recalling [17, Example 1.1]. Let us consider the line arrangement $\mathcal{L}\subset\mathbb{P}^{2}$ given by
$Q(x,y,z)=xyz(x+y-2z)(x-3y+z)(-5x+y+z)(x+y+z).$
This arrangement has $d=7$ with $t_{3}=3$ and $t_{2}=12$. One can check that $\mathcal{L}$ has type $2A$ with exponents $(4,4,4)$ and $\nu(\mathcal{L})=3$.
Moreover, a generic arrangement of $5$ lines described in Example 1.16 has $10$ nodes and exponents $(3,3,3,3)$. For this arrangement of type $2B$ we have equality in Theorem 1.25 (2).
2. Some preliminaries
With the notation from Introduction, we have the following result.
Theorem 2.1 ([7, Theorem 1.2]).
Let $C\,:f=0$ be a reduced plane curve of degree $d$ and $d_{1}={\rm mdr}(C)$. Then the following hold.
  1. (1)
    If $d_{1}<(d-1)/2$, then $\nu(C)=(d-1)^{2}-d_{1}(d-1-d_{1})-\tau(C)$.
  2. (2)
    If $d_{1}\geq(d-1)/2$, then
    $\nu(C)=\bigg{\lceil}\frac{3}{4}(d-1)^{2}\bigg{\rceil}-\tau(C).$
If we set $T=3(d-2)$, then the sequence $n(f)_{k}$ is symmetric with respect to the middle point $T/2$, that is one has
(2.1) $n(f)_{a}=n(f)_{b}$
for any integers $a,b$ satisfying $a+b=T$, see [29]. It was shown in [12, Corollary 4.3] that the graded $S$-module $N(f)$ satisfies a Lefschetz type property with respect to multiplication by generic linear forms. This implies, in particular, the inequalities:
(2.2) $0\leq n(f)_{0}\leq n(f)_{1}\leq\ldots\leq n(f)_{[T/2]}\geq n(f)_{[T/2]+1}\geq\ldots\geq n(f)_{T}\geq 0$
Moreover, for a $3$-syzygy curve $C$ of degree $d$ with exponents $(d_{1},d_{2},d_{3})$, we have the following formula for the initial degree of the graded module $N(f)$, see [17, Theorem 3.9],
(2.3) $\sigma(C)=\min\{k\ :\ N(f)_{k}\neq 0\}=3(d-1)-(d_{1}+d_{2}+d_{3}).$
2.2. Some key exact sequences
Let us come back to the setting of Theorem 1.4. To simplify the notation in the sequel we set
(2.4) $e_{1}=\deg f_{1}=\deg C_{1}\text{ and }e_{2}=\deg f_{2}=\deg C_{2}.$
Assume in addition that $C_{2}$ is smooth and all the points in $C_{1}\cap C_{2}$ are quasi-homogeneous singularities of $C=C_{1}\cup C_{2}$. Then there is an exact sequence for any integer $k$ given by
(2.5) $0\to E_{C_{1}}(k-e_{2})\stackrel{{\scriptstyle f_{2}}}{{\longrightarrow}}E_{C}(k)\to i_{2*}{\mathcal{O}}_{C_{2}}(-K_{C_{2}}-R+(k-1)D)\to 0,$
where $E_{C}=\widetilde{D_{0}(f)}$ (resp. $E_{C_{1}}=\widetilde{D_{0}(f_{1})}$ is the sheafification of the graded $S$-module $D_{0}(f)$ (resp. $D_{0}(f_{1})$), which is a rank two vector bundle on $\mathbb{P}^{2}$, $K_{C_{2}}$ is the canonical divisor on ${C_{2}}$, $R$ is the reduced scheme of $C_{1}\cap C_{2}$ and the divisor $D^{\prime}$ corresponds to the intersection of a line in $\mathbb{P}^{2}$ with the curve $C_{2}$, see [14, 28]. By taking the associated long cohomology exact sequence of the sequence (2.5), we get
(2.6) $0\to D_{0}(f_{1})_{k-e_{2}}\to D_{0}(f)_{k}\to H^{0}(C_{2},{\mathcal{O}}_{C_{2}}(D+(k-1)D^{\prime}))\to$
$\to N(f_{1})_{k-e_{2}+e_{1}-1}\to N(f)_{k+e_{1}+e_{2}-1}\to H^{1}(C_{2},{\mathcal{O}}_{C_{2}}(D+(k-1)D^{\prime}))\to$
$\to H^{2}(\mathbb{P}^{2},E_{C_{1}}(k-e_{2}))\to H^{2}(\mathbb{P}^{2},E_{C}(k))\to 0,$
with $D=-K_{C_{2}}-R$, and hence $\deg D=2-2g_{2}-r$, where $g_{C_{2}}$ is the genus of the smooth curve $C_{2}$ and $r$ is the number of points in the reduced scheme of $C_{1}\cap C_{2}$. Moreover, one has
(2.7) $\dim H^{2}(\mathbb{P}^{2},E_{C_{1}}(k-e_{2}))=\dim D_{0}(f_{1})_{n_{1}+n_{2}-(k-e_{2})-4}$
and
$\dim H^{2}(\mathbb{P}^{2},E_{C}(k))=\dim D_{0}(f)_{(n_{1}+n_{2}+1)-k-4},$
see [6, 29].
Remark 2.3.
If we identified $D_{0}(f)$ with the quotient $S$-module
$\overline{D}(f)=D(f)/SE,$
then the morphism $\theta$ is just given by the multiplication by $f_{1}$. Here $D(f)$ denotes the $S$-module of all derivations of $f$, that is
$D(f)=\{(a,b,c)\in S^{3}\ :\ af_{x}+bf_{y}+cf_{z}\in(f)\},$
where $(f)$ is the principal ideal generated by $f$, and $SE$ denotes the submodule in $D(f)$ generated by the Euler derivation $E=(x,y,z)$. The first nonzero morphism in (2.5) is multiplication by $f_{2}$ if we replace $D_{0}(f_{1})$ and $D_{0}(f)$ by $\overline{D}(f_{1})$ and $\overline{D}(f)$ respectively, see for instance [28, Theorem 1.6]. This fact is used in the proof of Theorems 1.19 and 1.24 below.
2.4. Exact sequences for line arrangements
Consider the triple of line arrangements $({\mathcal{A}}^{\prime},{\mathcal{A}},{\mathcal{A}}^{\prime\prime})$, where ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{L\}$, $L$ being a line in ${\mathcal{A}}$, and ${\mathcal{A}}^{\prime\prime}$ is the arrangement consisting of points on the line $L$ corresponding to the intersection set ${\mathcal{A}}^{\prime}\cap L$. By a change of coordinates we can take $L:x=0$ and set
${\mathcal{A}}^{\prime}:f^{\prime}=0,\ {\mathcal{A}}:f=0\text{ and }{\mathcal{A}}^{\prime\prime}:f^{\prime\prime}=0,$
for some polynomial $f^{\prime\prime}\in R=\mathbb{C}[y,z]$. Then there is an exact sequence of graded $S$-modules, sometimes called the Euler exact sequence of the triple $({\mathcal{A}}^{\prime},{\mathcal{A}},{\mathcal{A}}^{\prime\prime})$, given by
(2.8) $0\to D(f^{\prime})(-1)\to D(f)\to D(f^{\prime\prime}),$
where $D(f^{\prime})$, $D(f)$ and $D(f^{\prime\prime})$ are as in Remark 2.3 and in addition $D(f^{\prime\prime})$ is regarded as an $S$-module using the restriction of scalars given by the ring morphism $\psi:S\to R$, which replaces $x$ by 00 in any polynomial $h\in S$. Here the first morphism $D(f^{\prime})(-1)\to D(f)$ is the multiplication by $x$, and the second morphism is given by
$\rho:D(f)\to D(f^{\prime\prime}),\ \rho(a_{1},a_{2},a_{3})=(\psi(a_{2}),\psi(a_{3})),$
see [25, Proposition 4.45] and [27, Lemma 2.6] that covers a more general case when ${\mathcal{A}}^{\prime}$ is a conic-line arrangement, maybe with some quasi-homogeneity conditions. If we replace the modules $D(f^{\prime})$, $D(f)$ and $D(f^{\prime\prime})$ by the modules $\overline{D}(f^{\prime})$, $\overline{D}(f)$ and $\overline{D}(f^{\prime\prime})$, the sequence obtained from (2.8) is no longer exact. However, if we pass to the associated coherent sheaves on $\mathbb{P}^{2}$, we get the following short exact sequence
(2.9) $0\to E_{{\mathcal{A}}^{\prime}}(-1)\to E_{{\mathcal{A}}}\to i_{*}E_{{\mathcal{A}}^{\prime\prime}}\to 0,$
where we followed the notation from (2.5), see [26, Theorem 3.2]. Using this exact sequence we can give the following useful sufficient condition for the sequence (2.8) to be right-exact, i.e. for the morphism $\rho$ to be surjective.
Theorem 2.5.
With the above notation, let
$0\rightarrow\oplus_{i=1}^{m-2}S[-c^{\prime}_{i}]\rightarrow\oplus_{i=1}^{m}S[-d_{i}]\rightarrow D_{0}({\mathcal{A}}^{\prime})\rightarrow 0$
be a minimal free resolution of $D_{0}({\mathcal{A}}^{\prime})$. If $|{\mathcal{A}}^{\prime\prime}|>c_{i}^{\prime}-1$ for all $i$, then the Euler exact sequence
$0\rightarrow D({\mathcal{A}}^{\prime})[-1]\stackrel{{\scriptstyle\cdot x}}{{\rightarrow}}D({\mathcal{A}})\stackrel{{\scriptstyle\rho}}{{\rightarrow}}D({\mathcal{A}}^{\prime\prime})$
is right exact.
Proof.
Let $|{\mathcal{A}}^{\prime\prime}|=n$, and note that this implies $|{\mathcal{A}}^{\prime}|\geq n$. Then the $S$-module $D({\mathcal{A}}^{\prime\prime})$ has two generators, one of degree $1$ and another of degree $n-1$. The first element is the Euler derivation, which is in the image of $\rho$ always. It follows that
$D({\mathcal{A}})\stackrel{{\scriptstyle\rho}}{{\rightarrow}}D({\mathcal{A}}^{\prime\prime})$
is surjective if and only if the induced morphism
$\overline{D}({\mathcal{A}})\stackrel{{\scriptstyle\overline{\rho}}}{{\rightarrow}}\overline{D}({\mathcal{A}}^{\prime\prime})$
is surjective. Moreover, it suffices to show that $\overline{\rho}$ is surjective in degree $n-1$. Using the exact sequence (2.9), it suffices to show that $H^{1}(\mathbb{P}^{2},E_{{\mathcal{A}}^{\prime}}(n-2))=0$. By the minimal free resolution of $D_{0}({\mathcal{A}}^{\prime})$, we have
00 0 $=$ $H^{1}(\mathbb{P}^{2},\oplus\mathcal{O}(-d_{i}+n-2))\rightarrow H^{1}(\mathbb{P}^{2},E_{{\mathcal{A}}^{\prime}}(n-2))$
$\rightarrow$ $H^{2}(\oplus\mathcal{O}(-c^{\prime}_{i}+n-2))\simeq H^{0}(\mathbb{P}^{2},\oplus\mathcal{O}(c^{\prime}_{i}-n+2-3))=0.$
The last vanishing follows by our assumption, and the first one from the fact that $H^{1}(\mathbb{P}^{2},\mathcal{O}(k))=0$ for any integer $k$. This completes the proof.
The exact sequence in Theorem 2.5 plays a key role in our proof of Theorem 1.24. When we pass from line arrangements to conic-line arrangements, there are similar exact sequences according to [27], and they are discussed and used in the proof of Theorem 1.19 below.
3. The proofs of Theorems 1.4 and 1.7 and of Corollaries 1.9 and 1.11
3.1. The proof of Theorem 1.4
There is a natural injective $S$-linear morphism
$\theta:D_{0}(f_{1})_{k-e_{2}}\to D_{0}(f)_{k}$
for any integer $k$. If $\delta\in D_{0}(f_{1})_{k-e_{2}}$ is a syzygy, we can think about it as being a derivation killing $f_{1}$, and then we set
$\theta(\delta)=f_{2}\delta-\frac{\delta(f_{2})}{d}E,$
where $d=\deg C=e_{1}+e_{2}$ and $E$ denotes the Euler derivation, see [14, Theorem 5.1]. We denote by $d_{1},d_{2}$ (resp. by $d_{1}^{\prime},d_{2}^{\prime}$ ) the first two exponents of $C_{1}$ (resp. of $C$). By Proposition 1.3, we have
$d_{1}^{\prime}\leq d_{1}+e_{2}.$
If $d_{2}^{\prime}\leq d_{2}+e_{2}$, then by adding these two inequalities we get our claim. Suppose now that
$d_{2}^{\prime}>d_{2}+e_{2}.$
Let $r_{1}$ and $r_{2}$ be the first two generators in a minimal set of generators for $D_{0}(f_{1})$ and let $r_{1}^{\prime}$ be a generator of $D_{0}(f)$ of minimal degree $d_{1}^{\prime}$. Then $\theta(r_{1})\in D_{0}(f)_{d_{1}+e_{2}}$ and $\theta(r_{2})\in D_{0}(f)_{d_{2}+e_{2}}$ are non-zero derivations. It follows there are non-zero homogeneous polynomials $p,q\in S$ such that $\theta(r_{1})=pr_{1}^{\prime}$ and $\theta(r_{2})=qr_{1}^{\prime}$. This implies $q\theta(r_{1})-p\theta(r_{2})=\theta(qr_{1}-pr_{2})=0$, and hence by the injectivity of $\theta$ (see the beginning of the proof of [14, Theorem 5.1]) we get
$qr_{1}-pr_{2}=0.$
This is a contradiction since the quotient $S$-module $D_{0}(f_{1})/Sr_{1}$ is torsion free, as it follows, for instance, from the exact sequence in [18, Theorem 5.1].
3.2. The proof of Theorem 1.7 and of Corollaries 1.9 and 1.11
We use the exact sequence (2.6) above, where
$g_{2}=\operatorname{genus}(C_{2})=\frac{(e_{2}-1)(e_{2}-2)}{2}$
and $r=e_{1}e_{2}$. Moreover $D^{\prime}$ consists of $e_{2}$ points, and hence
$\deg(D+(k-1)D^{\prime})=2-(e_{2}-1)(e_{2}-2)-e_{1}e_{2}+(k-1)e_{2}=e_{2}(k-e_{1}-e_{2}+2).$
For $k<e_{1}+e_{2}-2$, we have $H^{0}(C_{2},{\mathcal{O}}_{C_{2}}(D+(k-1)D^{\prime}))=0$, and this yields an isomorphism
$D_{0}(f_{1})_{k-e_{2}}\to D_{0}(f)_{k}.$
It follows that all the generators of $D_{0}(f)$ of degree $k<e_{1}+e_{2}-2$ come from generators of $D_{0}(f_{1})$ of degree $k-e_{2}$. In this way we get two generators of degrees
$d_{1}^{\prime}=d_{1}+e_{2}\text{ and }d_{2}^{\prime}=d_{2}+e_{2}.$
A new generator, possibly coming from $H^{0}(C_{2},{\mathcal{O}}_{C_{2}}(D+(k-1)D^{\prime}))=\mathbb{C}$, arises when $k=e_{1}+e_{2}-2$. Since
$d_{2}^{\prime}=d_{2}+e_{2}\leq e_{1}+e_{2}-2$
by our assumption, we get that $d_{1}^{\prime}$ and $d_{2}^{\prime}$ are the first two exponents of the curve $C$, and hence $C$ has type
$d_{1}^{\prime}+d_{2}^{\prime}-(e_{1}+e_{2}-1)=(d_{1}+d_{2}-(e_{1}-1))+e_{2}=t_{1}+e_{2}$
as claimed.
The proof above implies that if the curve $C_{1}$ satisfies the condition $d_{1}\leq e_{1}-2$, then the resulting curve $C$ also satisfies this condition, since
$d_{2}^{\prime}=d_{2}+e_{2}\leq e_{1}+e_{2}-2.$
This completes the proof of Theorem 1.7.
Now we prove Corollary 1.9. The first claim follows from (1.3), since
$t({\mathcal{A}})=d_{1}+d_{2}-(d-1)\leq 2d-4-(d-1)=d-3.$
Moreover, (1.3) tells us that we can apply Theorem 1.7 to any line arrangement $C_{1}$. Starting with the integers $t$ and $d$ as in the second claim of Corollary 1.9, we choose $C_{1}$ to be a free line arrangement of $e_{1}=d-t\geq 3$ lines, for instance $x^{e_{1}}+y^{e_{1}}=0$. Then we add $t$ generic lines to this arrangement and we get by Theorem 1.7 a line arrangement ${\mathcal{A}}$ of degree $d$ and type $t$. This completes the proof of Corollary 1.9.
To prove Corollary 1.11, we first notice that our additional assumption on ${\mathcal{C}}$ implies
$t({\mathcal{C}})=d_{1}+d_{2}-\deg({\mathcal{C}})+1\leq 2d_{2}-\deg({\mathcal{C}})+1\leq\deg({\mathcal{C}})-3.$
Then, for the second claim, it is enough as above to construct free conic-line arrangements ${\mathcal{C}}_{1}$ for any $e_{1}=\deg({\mathcal{C}}_{1})=d-t\geq 3$, containing at least one conic and such that $d_{2}\leq\deg({\mathcal{C}}_{1})-2$. When $e_{1}=3$, we take ${\mathcal{C}}_{1}$ to be a smooth conic plus one of its tangents. It is known that then ${\mathcal{C}}_{1}$ is free with exponents $(1,1)$ and we are done.
For $e_{1}\geq 4$, we consider the arrangement
${\mathcal{C}}_{1}:f_{1}=x(xz+y^{2})(x^{e_{1}-3}+y^{e_{1}-3})=0.$
It is easy to check the following are two syzygies for $f_{1}$.
$r_{1}=(x^{2},xy,-e_{1}y^{2}-(e_{1}-1)xz)$
and
$r_{2}=(-(e_{1}-1)xy^{e_{1}-4},e_{1}x^{e_{1}-3}+y^{e_{1}-3},-2e_{1}x^{e_{1}-4}y+(e_{1}+1)y^{e_{1}-4}z).$
It follows that
$d_{1}\leq\deg(r_{1})=2\text{ and }d_{2}\leq\deg(r_{2})=e_{1}-3.$
since $r_{1}$ and $r_{2}$ are not multiple of strictly lower degree syzygy. Using Theorem 1.1, it follows that ${\mathcal{C}}_{1}$ is a free curve and the above inequalities are in fact equalities. In particular,
$d_{2}=\max\{2,e_{1}-3\}\leq e_{1}-2$
for $e_{1}\geq 4$. To complete the proof of Corollary 1.11 it remains to add $t$ generic lines to the arrangement ${\mathcal{C}}_{1}$ and get in this way the required arrangement ${\mathcal{C}}$ by Theorem 1.7.
Remark 3.3.
The curve ${\mathcal{C}}$ has a complicated singularity at the point $p=(0:0:1)$. Indeed, looking at $({\mathcal{C}},p)$ as the union of a $(e_{1}-2)$-ordinary point with the smooth germ at $p$ given by the conic $xz+y^{2}=0$ and using the known formula for the Milnor number of a union, see for instance [30, Theorem 6.5.1], one can show that the Milnor number $\mu({\mathcal{C}},p)$ is given by
$\mu({\mathcal{C}},p)=e_{1}^{2}-4e_{1}+6.$
On the other hand, using the formula for the total Tjurina number of a free curve, which can be obtained by setting $\nu(C)=0$ in Theorem 2.1 (1), we get
$\tau({\mathcal{C}})=(e_{1}-1)^{2}-2(e_{1}-3)=e_{1}^{2}-4e_{1}+7.$
Now the curve ${\mathcal{C}}$ has $e_{1}-3$ nodes $A_{1}$ as singularities except $p$, and hence
$\tau({\mathcal{C}},p)=\tau({\mathcal{C}})-(e_{1}-3)=e_{1}^{2}-5e_{1}+10.$
One can alternatively compute $\tau({\mathcal{C}},p)$ by using only local computations, based on [21, Proposition 3.2 and Theorem 5.1]. It follows that the invariant
$\epsilon({\mathcal{C}},p)=\mu({\mathcal{C}},p)-\tau({\mathcal{C}},p),$
which was introduced in [8] and which measure how far is $({\mathcal{C}},p)$ from being quasi-homogeneous, is given by
$\epsilon({\mathcal{C}},p)=e_{1}-4.$
Hence for all $e_{1}>4$ the singularity $({\mathcal{C}},p)$ is not quasi-homogeneous.
4. The proof of Propositions 1.13, 1.14 and 1.15
4.1. The proof of Proposition 1.13
To see that the number $m$ of syzygies can be only $3$ or $4$, and to get the corresponding values for the $\varepsilon_{j}$, use the equality (1.7). In case (1), the value for $\tau(C)$ follows from [17, Proposition 2.1] using that $d_{1}+d_{2}=d+1$, where $d=\deg C$. In case (2), we use (1.6) to get the following resolution of the Milnor algebra $M(f)$:
$0\to S(-d-d_{3})\oplus S(-d-d_{4})\to\oplus_{j=1}^{4}S(1-d-d_{j})\to S^{3}(1-d)\to S.$
This resolution implies that in this case one has
$\tau(C)=\binom{k+2}{2}-3\binom{k+3-d}{2}+\sum_{j=1}^{4}\binom{k+3-d-d_{j}}{2}-$
$-\binom{k+2-d-d_{3}}{2}-\binom{k+2-d-d_{4}}{2},$
for any large $k$. The last claim follows from the equality (1.7).
4.2. The proof of Propositions 1.14 and 1.15
We start with Proposition 1.14, and use Theorem 2.1 and Proposition 1.13. Note that the condition $d_{1}<d_{2}-2$ is exactly the condition $d_{1}<(d-1)/2$ which occurs in Theorem 2.1 (1). Hence, to prove the claim $(1)$ in Proposition 1.14, we use the formula for $\nu(C)$ given in Theorem 2.1 (1) and the formula for $\tau(C)$ given in Proposition 1.13, $(1)$. To prove the claims (2), (3) and (4), we use the formula for $\nu(C)$ given in Theorem 2.1 (2) and the formula for $\tau(C)$ given in Proposition 1.13 $(1)$. This completes the proof of Proposition 1.14.
The proof of Proposition 1.15 is analogous.
5. The proof of Theorem 1.18, Theorem 1.19, Proposition 1.22 and Theorem 1.24
5.1. The proof of Theorem 1.18
We use the exact sequence (2.6) above with $C_{1}={\mathcal{A}}^{\prime}:f_{1}=0$ and $C_{2}=L:f_{2}=0$. Hence $e_{1}=m_{1}+m_{2}=r$, $e_{2}=1$, $g_{2}=\operatorname{genus}(C_{2})=0$. Moreover $D^{\prime}$ is just a point, and hence
${\mathcal{O}}_{C_{2}}(D+(k-1)D^{\prime})={\mathcal{O}}_{L}(k+1-m_{1}-m_{2}).$
Let $(d_{1},d_{2},...,d_{m})$ be the exponents of the curve $C={\mathcal{A}}$.
Consider first the case $k<m_{1}+m_{2}-1=e_{1}-1$. Then $H^{0}(L,{\mathcal{O}}_{L}(k+1-m_{1}-m_{2}))=0$ and we get an isomorphism
$D_{0}(f_{1})_{k-1}\to D_{0}(f)_{k}.$
Note that both $k=m_{1}+1$ and $k=m_{2}+1$ satisfy the condition $k<m_{1}+m_{2}-1$, and hence we get that $d_{1}=m_{1}+1$ and $d_{2}=m_{2}+1$. It follows that $d_{1}+d_{2}=\deg{\mathcal{A}}+1$, and hence ${\mathcal{A}}$ has type $2$. To see what happens for the other exponents of ${\mathcal{A}}$, we look now at the case $k=e_{1}-1$. Since ${\mathcal{A}}^{\prime}$ is a $3$-syzygy curve, using (2.3) we get
$\sigma({\mathcal{A}}^{\prime})=\operatorname{indeg}N(f_{1})=3(e_{1}-1)-(m_{1}+m_{2}+m_{3})=2e_{1}-m_{3}-3.$
Let $T_{1}=3(e_{1}-2)$ and recall that $\dim N(f_{1})_{j}=\dim N(f_{1})_{T_{1}-j}$ for any integer $j$ by (2.1). It follows that $N(f_{1})_{j}=0$ for any $j>T_{1}-\sigma({\mathcal{A}}^{\prime})=e_{1}+m_{3}-3$. Since $m_{3}\leq e_{1}-2$ by (1.3), this yields an exact sequence
(5.1) $0\to D_{0}(f_{1})_{e_{1}-2}\to D_{0}(f)_{e_{1}-1}\to H^{0}(L,{\mathcal{O}}_{L})\to 0$
When $m_{3}=e_{1}-2$, then there are two new generators in $D_{0}(f)_{e_{1}-1}$, one corresponding to the generator of $D_{0}(f_{1})$ of degree $m_{3}$, and the second one corresponding to the generator of $H^{0}(L,{\mathcal{O}}_{L})=\mathbb{C}$. In this case
$(d_{1},d_{2},d_{3},d_{4})=(m_{1}+1,m_{2}+1,e_{1}-1,e_{1}-1)$
as claimed. When $m_{3}<e_{1}-2$, then the exact sequence (5.1) shows that there is exactly one new generator for $D_{0}(f)$ in degree $d_{3}=m_{3}+1<e_{1}-1$, and then a new generator in degree $d_{4}=e_{1}-1$ corresponding to the generator of $H^{0}(L,{\mathcal{O}}_{L})=\mathbb{C}$.
5.2. The proof of Theorem 1.19
This rather long proof is divided into two steps, to make its understanding easier for the reader.
Step 1. We use again the exact sequence (2.6) above with $C_{1}={\mathcal{C}}^{\prime}:f_{1}=0$ and $C_{2}=Q:f_{2}=0$. Hence now we have $e_{1}=m_{1}+m_{2}+1$, $r=2e_{1}$, $e_{2}=2$, $g_{2}=\operatorname{genus}(C_{2})=0$. Moreover, $D^{\prime}$ consists of $2$ points and hence
${\mathcal{O}}_{C_{2}}(D+(k-1)D^{\prime})={\mathcal{O}}_{L}(2k-2e_{1}).$
Let $(d_{1},d_{2},...,d_{m})$ be the exponents of the curve $C={\mathcal{C}}$.
Consider first the case $k<e_{1}$. Then $H^{0}(L,{\mathcal{O}}_{L}(2k-2e_{1}))=0$ and we get an isomorphism
$D_{0}(f_{1})_{k-1}\to D_{0}(f)_{k}.$
Note that both $k=m_{1}+2$ and $k=m_{2}+2$ satisfy the condition $k<m_{1}+m_{2}+1$, and hence we get that $d_{1}=m_{1}+2$ and $d_{2}=m_{2}+2$. It follows that $d_{1}+d_{2}=e_{1}+3=\deg{\mathcal{C}}+1$, and hence ${\mathcal{C}}$ has type $2$. To see what happens for the other exponents of ${\mathcal{C}}$, we look now at the case $k=e_{1}$. Since ${\mathcal{C}}^{\prime}$ is a free curve, one has $N(f_{1})=0$ and this yields an exact sequence
(5.2) $0\to D_{0}(f_{1})_{e_{1}-2}\to D_{0}(f)_{e_{1}}\to H^{0}(L,{\mathcal{O}}_{L})\to 0$
This exact sequence shows that there is exactly one new generator, call it ${\alpha}$, for $D_{0}(f)$ in degree $d_{3}=e_{1}$ corresponding to the generator of $H^{0}(L,{\mathcal{O}}_{L})=\mathbb{C}$.
Step 2. If the conic-line arrangement ${\mathcal{C}}$ has type $2A$, then the exponents of ${\mathcal{C}}$ are $(d_{1},d_{2},d_{3})=(m_{1}+2,m_{2}+2,m_{1}+m_{2}+1)$, and our result is proved. To show that this is the case, we have to show that there is no need of an additional generator for $D_{0}(f)$ beyond ${\alpha}$.
To do this, we recall some key results from [27]. A triple $({\mathcal{C}}^{\prime},{\mathcal{C}},{\mathcal{C}}^{\prime\prime})$ of CL (short for conic-line) arrangements, where ${\mathcal{C}}^{\prime}={\mathcal{C}}\setminus\{Q\}$ with $Q\in{\mathcal{C}}$ a smooth conic and ${\mathcal{C}}^{\prime\prime}$ the 0-dimensional arrangement on $Q$ given by the reduced intersection ${\mathcal{C}}^{\prime}\cap Q$, is called quasi-homogeneous if all the singular points of ${\mathcal{C}}^{\prime}$ and of ${\mathcal{C}}$ situated on $Q$ are quasi-homogeneous. The authors in [27] require that all the singular points of ${\mathcal{C}}^{\prime}$ and of ${\mathcal{C}}$ are quasi-homogeneous, but this stronger form is not necessary. Indeed, as noted in [14] and in subsection 2.2 above, for a similar condition in [28], this assumption is used in [27] only to compute the difference
$\tau({\mathcal{C}})-\tau({\mathcal{C}}^{\prime})=\deg J_{f}-\deg J_{f^{\prime}},$
where we set
${\mathcal{C}}^{\prime}:f^{\prime}=0\text{ and }{\mathcal{C}}:f=0,$
see for instance the proof of [27, Proposition 2.8 and Proposition3.7]. To do this computation, our weaker assumption is sufficient.
In [27, Proposition 3.5], the following exact sequence of $S$-modules is constructed
(5.3) $0\to D(f^{\prime})(-2)\to D(f)\to D(f^{\prime\prime})_{\psi},$
where $D(f)$, $D(f^{\prime})$ are as in Remark 2.3. To explain the equation ${\mathcal{C}}^{\prime\prime}:f^{\prime\prime}=0$ and the $S$-module $D(f^{\prime\prime})_{\psi}$, note that by a change of coordinates on $\mathbb{P}^{2}$ we can assume that $Q:q=y^{2}-xz=0$. Then we have an isomorphism
$v:\mathbb{P}^{1}\to Q\text{ given by }v(s,t)=(s^{2},st,t^{2}).$
Using this isomorphism, we can regard the 0-dimensional arrangement ${\mathcal{C}}^{\prime\prime}$ as a 0-dimensional arrangement in $\mathbb{P}^{1}$ and hence it is defined by an equation $f^{\prime\prime}=0$, where $f^{\prime\prime}\in R=\mathbb{C}[s,t]$ is some homogeneous polynomial. Hence the graded $R$-module $D(f^{\prime\prime})$ is well defined. Moreover, the map $v$ above induces a ring morphism
$\psi=v^{*}:S\to R\text{ given by }v^{*}(g)=g(v(s,t)),$
and the $S$-module $D(f^{\prime\prime})_{\psi}$ is just the module obtained from the $R$-module $D(f^{\prime\prime})$ by restriction of scalars using $\psi$. Next we consider the morphisms in the exact sequence (5.3). The first morphism $D(f^{\prime})(-2)\to D(f)$ is given by the multiplication by $q$, the equation of $Q$, and hence it preserves the homogeneous components, that is
$q(D(f^{\prime})(-2)_{k})=q(D(f^{\prime})_{k-2})\subset D(f)_{k}.$
The second morphism, call it $\rho:D(f)\to D(f^{\prime\prime})_{\psi}$, is much more subtle. If $(a_{1},a_{2},a_{3})\in D(f)$, then it is shown that
$\psi(a_{1})=sQ_{1},\ \psi(a_{2})=\frac{tQ_{1}+sQ_{2}}{2}\text{ and }\psi(a_{3})=tQ_{2}$
for some polynomials $Q_{1},Q_{2}\in R$, see [27] for details. With this notation, one sets
$\rho(a_{1},a_{2},a_{3})=(Q_{1},Q_{2})$
and hence
$\rho(D(f)_{k})\subset D(f^{\prime\prime})_{2k-1}.$
After these preliminaries, we can continue the proof in Step 2. In our situation, ${\mathcal{C}}^{\prime\prime}$ consists of $2e_{1}$ points, hence $f^{\prime\prime}$ is a homogeneous polynomial of degree $2e_{1}$ without multiple factors. The derivation ${\alpha}\in D_{0}(f)_{e_{1}}$ that was constructed at the end of Step 1 can be regarded as sitting in $\overline{D}(f)_{e_{1}}$ using the identification from Remark 2.3. Then any representative ${\alpha}_{0}\in D(f)_{e_{1}}$ of ${\alpha}$ is clearly not in the image of the morphism $D(f^{\prime})(-2)\to D(f)$ and hence $\rho({\alpha}_{0})\neq 0$. If we assume ${\alpha}_{0}=(a_{1},a_{2},a_{3})$, then
(5.4) $za_{1}-xa_{3}\neq 0.$
Indeed, $za_{1}=xa_{3}$ implies that there is $h\in S$ such that $a_{1}=xh$ and $a_{3}=zh$. But this implies that
${\alpha}_{1}={\alpha}_{0}-hE=(0,a_{2}^{\prime},0)$
is a new representative of ${\alpha}$ and clearly $\rho({\alpha}_{1})=0$, a contradiction.
If ${\mathcal{C}}$ has type $2B$, it means that there is an additional generator of degree $d_{4}>d_{3}=e_{1}$. Using (1.4), we get $d_{4}\leq(e_{1}+2)-1$, and hence $d_{4}=e_{1}+1$. On the other hand, the exact sequence (5.2) where $e_{1}$ is replaced by $e_{1}+1$ shows that $\overline{D}_{0}(f)_{e_{1}+1}$ modulo the image of $\overline{D}_{0}(f_{1})_{e_{1}-1}$ is 2 dimensional. It follows that we need a new generator of degree $e_{1}+1$ if and only the vector space $V$ spanned by $x{\alpha}_{0},y{\alpha}_{0}$ and $z{\alpha}_{0}$ in $\overline{D}_{0}(f)_{e_{1}+1}$ modulo the image of $\overline{D}(f_{1})_{e_{1}-1}$ is 1-dimensional. Note that $\rho(E)=E^{\prime\prime}$, the Euler derivation corresponding to the ring $R$, and hence there is an induced morphism
$\overline{\rho}:\overline{D}(f)\to\overline{D}(f^{\prime\prime})_{\psi}.$
The element
$x{\alpha}_{0}=(xa_{1},xa_{2},xa_{3})$
is equivalent modulo $SE$ to
${\beta}=x{\alpha}_{0}-a_{1}E=(0,xa_{2}-ya_{1},xa_{3}-za_{1}).$
Since $xa_{3}-za_{1}=s^{2}tQ_{2}-t^{2}sQ_{1}=st(sQ_{2}-tQ_{1})$, it follows that
(5.5) ${\beta}_{1}=\overline{\rho}(x{\alpha})=\overline{\rho}({\beta})=(0,s(sQ_{2}-tQ_{1}))\in\overline{D}(f^{\prime\prime})_{\psi}.$
A similar computation shows that
(5.6) ${\beta}_{2}=\overline{\rho}(z{\alpha})=(-t(sQ_{2}-tQ_{1}),0)\in\overline{D}(f^{\prime\prime})_{\psi}.$
The elements ${\beta}_{1}$ and ${\beta}_{2}$ are linearly independent in $\overline{D}(f^{\prime\prime})_{\psi}$. Indeed, no linear combination $c_{1}{\beta}_{1}+c_{2}{\beta}_{2}$ with $c_{1},c_{2}\in\mathbb{C}$ can have the form $hE^{\prime\prime}=(sh,th)$ for some $h\in R$. Indeed
$c_{1}{\beta}_{1}+c_{2}{\beta}_{2}=(-c_{2}t(sQ_{2}-tQ_{1}),c_{1}s(sQ_{2}-tQ_{1}))=(sh,th)$
would imply $(c_{1}s^{2}+c_{2}t^{2})(sQ_{2}-tQ_{1})=0$, which is a contradiction by (5.4). This shows that the classes of $x{\alpha}_{0}$ and of $z{\alpha}_{0}$ in $\overline{D}(f)$ are linearly independent as well. Therefore $\dim V\geq 2$ and this completes the proof of Theorem 1.19.
5.3. The proof of Proposition 1.22
The claims for $n_{1}=1$ and $n_{1}=2$ follow from [10, Proposition 4.7 (4)]. Consider now the case $n_{1}\geq 3$. Let ${\mathcal{B}}:g=0$ be the line arrangement obtained from ${\mathcal{A}}^{\prime}$ by adding the line $L^{\prime}$ determined by the points $a_{1}$ and $a_{2}$. Then ${\mathcal{B}}$ is a supersolvable line arrangement, and $a_{1}$ is a modular point of multiplicity $m_{1}=n_{1}+1$. Using [16, Proposition 3.2], it follows that
$mdr(g)=\min\{m_{1}-1,(n_{1}+n_{2}+1)-m_{1}\}=n_{1}.$
Since any supersolvable arrangement is free, it follows that ${\mathcal{B}}$ is a free arrangement with exponents $(n_{1},n_{2})$. Next, we apply [8, Theorem 1.3] to the curve $C={\mathcal{A}}^{\prime}$ and the line $L=L^{\prime}$. In our case, $r=|{\mathcal{A}}^{\prime}\cap L^{\prime}|=2$ and $\epsilon:=\epsilon({\mathcal{A}},L^{\prime})=0$ since all the singularities of line arrangements are quasi-homogeneous. It follows that the curve ${\mathcal{A}}^{\prime}$ is plus-one generated with exponents $(n_{1},n_{2},n_{1}+n_{2}-2)$. Indeed, we have to use the Case $(3)$ in [8, Theorem 1.3], where $d=\deg(C)=\deg({\mathcal{A}}^{\prime})=n_{1}+n_{2}$. In the notation of [8, Theorem 1.3], we have $d_{1}^{\prime}=n_{1}$ and hence $d_{1}\geq d_{1}^{\prime}-1=n_{1}-1$. It follows that
$d_{1}+1-\epsilon\geq n_{1}\geq 3>r,$
and hence by the last claim in [8, Theorem 1.3] the curve $C$ is not free. Therefore we are in the case (3) of [8, Theorem 1.3], and this yields our claim.
5.4. The proof of Theorem 1.24
This proof is similar to the proof of Theorem 1.19, so we divide it into two steps.
Step 1. We use the exact sequence (2.6) above with $C_{1}={\mathcal{A}}^{\prime}:f_{1}=0$ and $C_{2}=L:f_{2}=0$. Hence $e_{1}=n_{1}+n_{2}$, $e_{2}=1$, $g_{2}=\operatorname{genus}(C_{2})=0$, $r=n_{1}+n_{2}-1$. Moreover, $D^{\prime}$ is just a point and we have
${\mathcal{O}}_{C_{2}}(D+(k-1)D^{\prime})={\mathcal{O}}_{L}(k+2-n_{1}-n_{2}).$
Let $(d_{1},d_{2},...,d_{m})$ be the exponents of the curve $C={\mathcal{A}}$.
Consider first the case $k<n_{1}+n_{2}-2=e_{1}-2$. Then $H^{0}(L,{\mathcal{O}}_{L}(k+2-d_{1}-d_{2}))=0$ and we get an isomorphism
$D_{0}(f_{1})_{k-1}\to D_{0}(f)_{k}.$
Note that for $n_{1}\geq 4$, both $k=n_{1}+1$ and $k=n_{2}+1$ satisfy the condition $k<n_{1}+n_{2}-2$, and hence we get that $d_{1}=n_{1}+1$ and $d_{2}=n_{2}+1$. To see what happens for $n_{1}=3$, we look now at the case $k=e_{1}-2$. Since ${\mathcal{A}}^{\prime}$ is a 3-syzygy curve, using (2.3) we get
$\sigma({\mathcal{A}}^{\prime})=\operatorname{indeg}N(f_{1})=3(e_{1}-1)-(n_{1}+n_{2}+n_{1}+n_{2}-2)=e_{1}-1.$
Let $T_{1}=3(e_{1}-2)$ and recall that $\dim N(f_{1})_{j}=\dim N(f_{1})_{T_{1}-j}$ for any integer $j$ by (2.1). It follows that $N(f_{1})_{j}=0$ for any $j>T_{1}-\sigma({\mathcal{A}}^{\prime})=2e_{1}-5$. In particular, this yields an exact sequence
(5.7) $0\to D_{0}(f_{1})_{e_{1}-3}\to D_{0}(f)_{e_{1}-2}\to H^{0}(L,{\mathcal{O}}_{L})\to 0$
Consider first the case $n_{1}=3<n_{2}$, when the sequence becomes
$0\to D_{0}(f_{1})_{n_{2}}\to D_{0}(f)_{n_{2}+1}\to H^{0}(L,{\mathcal{O}}_{L})\to 0$
This shows that there are two new generators in $D_{0}(f)_{n_{2}+1}$, one corresponding to the generator of $D_{0}(f_{1})$ of degree $n_{2}$ and the other one, call it ${\alpha}$, corresponding to the generator of $H^{0}(L,{\mathcal{O}}_{L})=\mathbb{C}$. This implies that $d_{2}=d_{3}=n_{2}+1$ in this case. When $n_{1}=3=n_{2}$, then there are three new generators in $D_{0}(f)_{n_{2}+1}$, two corresponding to the two generators of $D_{0}(f_{1})$ of degree $n_{1}=n_{2}$, and the third one, call it again ${\alpha}$, corresponding to the generator of $H^{0}(L,{\mathcal{O}}_{L})=\mathbb{C}$. In this case $(d_{1},d_{2},d_{3})=(4,4,4)$ as claimed. It follows that in all cases ${\mathcal{A}}$ has type $2$. Now consider the case $n_{1}\geq 4$ and study the exponents $d_{3}$ and $d_{4}$. The exact sequence (5.7) shows that there is exactly one new generator for $D_{0}(f)$ in degree $d_{3}=e_{1}-2=n_{1}+n_{2}-2$, which we denote by ${\alpha}$, corresponding to the generator of $H^{0}(L,{\mathcal{O}}_{L})=\mathbb{C}$.
Step 2. If ${\mathcal{A}}$ is a $3$-syzygy curve, then Theorem 1.24 is proved. We show now that this is indeed the case.
The arrangement ${\mathcal{A}}^{\prime\prime}$ consists of $n=e_{1}-1=n_{1}+n_{2}$ points, using the formulas (1.6) and (1.7), we see that $m=3$, $\varepsilon_{1}=1$ and $c_{1}=2(n_{1}+n_{2}-1).$ Hence the integer $c^{\prime}_{1}$ from Theorem 2.5 is given by
$c_{1}^{\prime}=c_{1}-(e_{1}-1)=n_{1}+n_{2}-1.$
It follows that we can apply Theorem 2.5 and get a surjective morphism $\rho$. Now our claim that $D_{0}(f)$ can be generated by 3 elements or, equivalently, that $D(f)$ is generated by 4 elements, follows from the next basic fact in commutative algebra. If we have an exact sequence of $S$-modules
$0\to M^{\prime}\to M\to M^{\prime\prime}\to 0,$
$A$ a generating set for $M^{\prime}$ and a set $B\subset M$ which maps onto a generating set for $M^{\prime\prime}$, then the image of $A$ under $M^{\prime}\to M$ union $B$ gives a generating set for $M$. We apply this to $M^{\prime}=D(f^{\prime})$, $M=D(f)$ and $M^{\prime\prime}=D(f^{\prime\prime})$, and $A$ consisting of $E^{\prime}=E$ the Euler derivation of ${\mathcal{A}}^{\prime}$ and of ${\mathcal{A}}$, plus the two generators of $D(f^{\prime})_{0}$ and $B$ consisting of $E$ and an element ${\alpha}\in D(f)$ such that $\rho({\alpha})$ is the generator of $D(f^{\prime\prime})$ of degree $n-1$. The element $xE$ is discarded from the union since we have $E$ in this union.
6. The proof of Theorem 1.25
Let us recall that by [13, Theorem 2.1] if $C$ is a reduced curve of degree $d$ in $\mathbb{P}^{2}$ having only quasi-homogeneous singularities, then
$d_{1}\geq\alpha_{C}\cdot d-2,$
where $\alpha_{C}$ denotes the Arnold exponent of $C$. Let us recall that the Arnold exponent of a given curve $C$ is the minimum over all log canonical thresholds of singular points of $C$. Coming back to our setting, since for line arrangements all singular points are quasi-homogeneous, we can use the aforementioned result, but first we need to find $\alpha_{\mathcal{L}}$. Using [3, Theorem 1.3], we see that
$\alpha_{\mathcal{L}}=\frac{2}{m(\mathcal{L})},$
and we get
$d_{1}\geq\frac{2d}{m(\mathcal{L})}-2.$
Since $d_{1}\leq(d+1)/2$, we finally obtain
$m(\mathcal{L})\geq\bigg{\lceil}\dfrac{4d}{d+5}\bigg{\rceil}.$
To prove the claim in (1), we use Proposition 1.13 (1). Since $d^{2}=d_{1}^{2}+d_{2}^{2}+2d_{1}d_{2}-2d_{1}-2d_{2}+1$, we get
$\tau(\mathcal{L})=d^{2}-d_{1}d_{2}-2d_{3}-1.$
Recall that
$\tau(\mathcal{L})=\sum_{r\geq 2}(r-1)^{2}t_{r}=\sum_{r\geq 2}(r^{2}-2r+1)t_{r}.$
Since $d^{2}-d=\sum_{r\geq 2}(r^{2}-r)t_{r}$, we have
$\tau(\mathcal{L})=d^{2}-d+\sum_{r\geq 2}(-r+1)t_{r}.$
This gives us
$d^{2}-d+\sum_{r\geq 2}(-r+1)t_{r}=d^{2}-d_{1}d_{2}-2d_{3}-1,$
and finally
$\sum_{r\geq 2}(r-1)t_{r}=d_{1}d_{2}-d+2d_{3}+1=d_{1}d_{2}-d_{1}-d_{2}+2d_{3}+2.$
The claim follows now from the fact that $d_{1}\leq d_{2}\leq d_{3}$.
To prove the claim in (2), we use Proposition 1.13 (2). As above we get
$\tau(\mathcal{L})=d^{2}-d_{1}d_{2}-d_{3}-d_{4}.$
This yields
$d^{2}-d+\sum_{r\geq 2}(-r+1)t_{r}=d^{2}-d_{1}d_{2}-d_{3}-d_{4},$
and hence
$\sum_{r\geq 2}(r-1)t_{r}=d_{1}d_{2}+d_{3}+d_{4}-d_{1}-d_{2}+1.$
The claim (2) follows, since $d_{1}\leq d_{2}\leq d_{3}\leq d_{4}$.
Acknowledgement
Takuro Abe is partially supported by JSPS KAKENHI Grant Numbers JP23K17298 and JP23K20788.
Piotr Pokora is supported by the National Science Centre (Poland) Sonata Bis Grant 2023/50/E/ST1/00025. For the purpose of Open Access, the author has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.